Operations Report – February 12th

MDRS Supplemental Operations Report 12-FEB-2020 (Updated)

Name of person filing report: Shannon Rupert

Reason for Report: Routine Update/Return to campus

Non-nominal systems: Generator

Action is taken for non-nominal systems: Currently on a rental generator.

We determined that the generating unit has failed in the generator.
Because it has been replaced before (and it about half the price of a
new generator) and the engine has over 18K hours, we will be replacing the generator. The plan is to keep the rental for another month while finding a 10K replacement generator that is extended use and is compatible with the power system’s autostart. The power system itself is nominal.

Generator check, note if oil and coolant added: Checked levels on the
rental, all are nominal at this time.

ScienceDome AC unit, on and operational: Still not operational.

Solar— running nominally, solar throughout the day and generator
backup usually on 12 hours a night

Diesel Reading – not checked
Propane Reading – not checked
Ethanol Free Gasoline – not checked

Water (director’s tank) – not checked
Water (static tank) – not checked
Water (GreenHab) – not checked
Water (loft) – not checked
Static to Loft Pump used – n/a
Water Meter: n/a

Toilet tank emptied: n/a

Deimos rover used: Yes. Repairs completed. This rover has new
batteries which need to be used 30 times before they are fully
charged. This rover can now be used by crews as long as:

1. Rover is not taken out as the only rover
2. Rovers still return to station when the first one reaches 60 %
3. Keep an eye on the charge throughout the EVA. If you see a quick drop in the charge (say from 70 to 40 % in a few seconds) come home immediately.
4. You can take the rover out with just one person in each rover
while this rover is fully charging its batteries.

Beginning charge:
Ending charge:
Currently charging: yes

Sojourner rover used: nominal
Beginning charge:
Ending charge:
Currently charging: yes

Spirit rover used: nominal
Beginning charge:
Ending charge:

Currently charging: yes. This rover now has the metal parts that
ripped the spacesuits covered. It is a good plan and we will cover
the others as soon as possible.

Opportunity rover used: Still has the brake issue and should not be
used until I work things out with the repair shop who failed to make
the repair. Other than the brake issue, it is nominal.

Beginning charge:
Ending charge:
Currently charging: yes

Curiosity rover used: nominal.
Beginning charge:
Ending charge:
Currently charging: yes

Notes and action took on rovers: I changed the power cord that was
being used to charge Opportunity. It was missing its grounding prong.
In addition, I secured both charging ports at the RAM so the plug is

Monthly check and fill of rover batteries: Not done at this time.
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): 350.1 is the only ATV
operating nominally. The 300 is still in town. 350.2 has a safety
issue with the damaged handlebars and currently does not have a
battery. 350.2 and the Honda have low batteries

Reason for use: Started and checked for operations
Oil Added? No
ATV Fuel Used: 0 Gals
# Hours the ATVs were Used: n/a

Notes and action took on ATVs: I have the jumper cables from the Hab and will work on them tomorrow. I’m also going to try to adjust the handlebars on the damaged rover.

HabCar running? If used, why, where? Yes, used for trips to town. It
has the wrong battery in it (it’s the battery is currently on the
generator) and it is filthy. Needs new jumper cables and its plastic
shovel for mud

CrewCar running, If used why, where? Currently in GJ. Needs to be
returned to complete repairs that were not completed last time it was
in the shop (Atila, please confirm)

Campus-wide inspection, if an action is taken, what and why? Mostly nominal.
Small diesel spill at the tank has gotten larger and will need
remediation. The leak has been stopped. Rockpile in front of Hab has
been removed. Looks great. Someone has created a new and incorrect route to deliver water than goes between the Hab and the observatory and solar array. In the future, the correct route needs to be used.

Evidence of rodents, where and action is taken: Two rodents caught in Hab and released. Crews need to make sure all food is in cupboards and the kitchen is clean every night. There is evidence that a rodent was in my trailer, but apparently they have moved out.

Summary of internet: All three accounts nominal. Suggest replacement of wifi box in Hab (I can never remember which is the router and which is the modem)

Summary of suits and radios: Four yellow radios were added to the
crew supply. Two were from the ScienceDome originally so only two new ones were added. Radios are the most replaced item on campus. Did not check the suits.

Summary of Hab operations: Nothing to report at this time as the crew is in sim

Summary of Outpost operations: The discharge plumbing to the intern
trailer has been installed. However, the replacement of dirt is still a
work in progress and needs to be completed.

Summary of GreenHab operations: New seed bank can and additional seeds
were added to the GreenHab. Nothing else to report as the crew is in
sim at this time.

Summary of ScienceDome operations: Dual split still not completed.
The incubator is non-functional due to a failure in its digital chip.
Crews will need to use the oven for an incubator. Additional soil
needs to be banked up and over the cement foundation outside to
protect the interior floor from water damage.

Summary of RAM operations: The wiring issue at the charging stations has been completed. There is new damage to the top of the outer airlock door and no one seems to know how it happened.

Summary of any observatory issues: A cord was replaced by Atila and
Peter and both observatories are in good shape.

Summary of health and safety issues: The stove was installed with part of the changeover from natural gas to propane not completely done.

The crew identified the problem and it has been corrected. Stove
operations are now nominal.

List and summary of projects: Now that I am back on campus, I will be
working with Marc and Mike from Crew 216 to rework the MDRS map.
I’m excited about this project!

Questions, concerns, supplies needed and requests: Nothing at this time.

Copyright © The Mars Society. All rights reserved. | Main Site